OpenStudy (anonymous):

Diana invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of her investment after x years. f(x)-400(1.01)^x

OpenStudy (anonymous):

I need to find a way to do it please

OpenStudy (anonymous):

1.01 dollars per year 2.47 dollars per year 4.12 dollars per year 12.36 dollars per year

OpenStudy (wolf1728):

I'm not exactly sure of your question. f(x)-400(1.01)^x Is that total = 400*(1.01)^years ? 400*1.01^1 = $404.00 and are those 4 amounts the multiple choices?

OpenStudy (anonymous):

oops sorry, I meant to put f(x)=400(1.01)^x And yes the amounts are multiple choice

OpenStudy (wolf1728):

So, after 1 year the amount is 404.00 Looks as if $4.12 is the closest.

OpenStudy (anonymous):

Alright thank you

OpenStudy (anonymous):

Well actually this asks for rate of change from the second year to the fifth

OpenStudy (anonymous):

Calculate the value at the end of the firth year, A5, and the end of the second year, A2, then the rate of change will be (A5-A2)/(5-2) per year